Suppose the inverse market demand curve for widgets is given by p = 12 – 2Q, and the market is characterized by Stackelberg duopoly. Both firms have marginal costs of 2 and fixed costs of 0. What is the equilibrium price in the market? Your answer should be rounded to the first decimal place (e.g. 123.4).  
Answer #1

Under Stackelberg duopoly oue Hum mouer first and second will follow it so we are assuming that firm , mover. first and firmTi = __102) - 293, 2에 톰 -월) = 102 - 29로 594 오로 a . 이외 - put 5 . 오 5 - 231 diti - o da, 522) 20 오2 들길들) ! AAN 여 21t2 = 틀 +득 뿜